# MagicTime: Time-lapse Video Generation Models as Metamorphic Simulators

### v1

@@ 3.2 ChronoMagic Dataset
where $i$ represents the position of the frame. Ultimately, we segment the video into distinct sections using the identified transition points. Due to possible detection errors in OpenCV and CLIP, the video may still contain transitions that are challenging to identify after processing. In the subsequent stage, we enhance video quality by manually correcting incorrectly cropped segments and then screening clean segments. As illustrated in Fig. 2a, the process results in a high-quality video dataset with fewer transitions.
Multi-view Text Fusion    After preprocessing, it is essential to annotate the video with appropriate captions. A straightforward approach is to input keyframes into a multimodal model for video description. However, the absence of substantial intermediate information may compromise the quality of the annotations. The ChronoMagic dataset’s language modality includes multiview text, such as titles, hashtags, keyframe captions, and video captions, so we leverage these to enhance labeling accuracy. Fig. 2b shows the detailed text generation and fusion process. In videos, titles underscore main themes and actions, while hashtags highlight focus and dynamics. Theoretically, high-quality captions could be generated solely from titles and hashtags. Nevertheless, exaggerated titles and hashtags on the internet can diminish caption quality. To counteract this, we employ in-context Learning and Chain-of-Thought (ICL, ) to guide GPT-4V (gpt4, ; lyu2023gpt, ) in generating keyframe captions using titles and hashtags, thereby reducing model bias and capturing spatial video information. Moreover, temporal video information is crucial. Thus, we use GPT-4V to derive a comprehensive representation of the entire video from keyframe captions, resulting in the final video captions. Additionally, at this stage, GPT-4V evaluates whether the video is a time-lapse and then forms a closed data loop as described in Section 3.2.
3.3 MagicTime Model
MagicAdapter    To transform the pre-trained T2V model to generate metamorphic videos, we decouple the training process into two stages: spatial training and temporal training, refer to Fig. 3.
Internet video data often contains non-removable watermarks, introducing irrelevant noise and data bias. Video generation models consist of spatial and temporal layers—the former focuses on learning visually relevant content, while the latter builds inter-frame continuity. If we can separate temporal and spatial training, allowing the spatial module to concentrate more on the subject rather than the watermark, this will facilitate the temporal layer in learning the motion patterns of metamorphic videos. To accomplish this, in the initial training phase, we remove the temporal layer from the pre-trained model and integrate the MagicAdapter-S into the spatial layer, inspired by QLoRA Adapter (Qlora, ). Subsequently, we freeze the remaining model parameters and employ keyframe-text pairs from the ChronoMagic dataset for training. Formally, let input $X$ to a single spatial layer $\mathcal{F}_{sap}$ and MagicAdapter-S layer $\mathcal{R}_{ada}$ be in the form $\mathbb{R}^{\text{B}\times\text{C}\times\text{H}\times\text{W}}$. The output of this layer is:
$\mathcal{F}_{out}(X)=\mathcal{F}_{sap}(X)+\alpha\times\mathcal{R}_{ada}(X),$ (8)
where $\mathcal{F}_{out}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{H}\times\text{W}}$, and $\alpha$ is a scaling coefficient used to regulate the scope of the adapter. $B$, $C$, $F$, $H$, and $W$ respectively represent the batch size, channel, number of frames, height, and width of the video. Ensuring high temporal consistency is crucial for video generation models. While general videos require consistency within localized segments, metamorphic videos necessitate consistency throughout the entire video, capturing a completely metamorphic process. Training alongside general videos could hinder the model’s comprehension of the new world phenomena. To mitigate this, we separate the training of general and metamorphic videos, thereby maintaining the model’s ability to generate general videos. Specifically, in the second training stage, we reintroduce the pre-trained temporal layer from the first stage, freeze all parameters, and propose a MagicAdapter-T into the temporal layer. Subsequently, we freeze other parameters and conduct training using video-text pairs from the ChronoMagic dataset. If the input $X$ to a single temporal layer $\mathcal{F}_{temp}$ and MagicAdapter-T layer $\mathcal{M}_{ada}$ is in the form of $\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$, then the output of this layer is given by:
$\mathcal{F}_{out}(X)=\mathcal{F}_{temp}(X)+\beta\times\mathcal{M}_{ada}(X),$ (9)
where $\mathcal{F}_{out}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$, $\beta$ is a scaling coefficient, and $B$, $C$, $F$, $H$, and $W$ respectively represent the batch size, channel, number of frames, height, and width of the video. Based on large-scale video-text data (webvid, ), the pre-trained model incorporates general motion capabilities (animatediff, ) through its temporal layer. MagicAdapter-T further enhances the model’s ability to generate metamorphic videos by integrating physical knowledge and predicting metamorphic motion.
@@ 3.3 MagicTime Model
Table 2: Comparison of statistics among Sky Time-lapse (SkyTimelapse, ), Time-lapse-D (TimeLapse-D, ), and the proposed metamorphic ChronoMagic dataset. ChronoMagic is the only high-quality dataset for text-to-video generation tasks.
ChronoMagic Sky Time-lapse (SkyTimelapse, ) Time-lapse-D (TimeLapse-D, )
@@
Resolution 1280 $\times$ 720 640 $\times$ 360 1920 $\times$ 1080
Dynamic Frames Extraction    Open-source algorithms, whether for video generation (360DVD, ; Magic-Me, ) or multimodal models (Video-llava, ; Languagebind, ), typically sample a random continuous sequence of $N$ frames for training. However, this approach captures only a brief segment of the full video, leading to limitations such as monotonous actions and restricted variation in amplitude. This is not conducive to extracting training frames for metamorphic videos. By analyzing ChronoMagic’s characteristics, we opt to sample $N$ frames evenly from a video, capitalizing on the nature of metamorphic videos that encompass the entire process of object metamorphosis. This ensures that the training data exhibit metamorphic properties, rather than being confined to general ones, thereby enhancing the generative models’ ability to comprehend physical phenomena. Nonetheless, uniform sampling from all videos might compromise the model’s original capabilities, resulting in generated videos that lack the action continuity of general videos. Consequently, we retain a portion of general videos and dynamically select the sampling strategy based on the results of cascade preprocessing:
$\mathcal{S}^{*}=\begin{cases}\text{Uniform, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}>\delta\\ \text{Random, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}\leq\delta\end{cases}$ (10)
Magic Text-Encoder    Textual descriptions corresponding to metamorphic videos typically contain more temporal and state information than those for general videos. Consequently, using a text encoder trained exclusively on general videos is deemed insufficient for this purpose. To overcome this limitation, we introduce the Magic Text-Encoder, which is specifically designed to encode metamorphic prompts while retaining its ability to process general prompts. We select CLIP as our initial text encoder and incorporate a Low-rank Adapter to encode metamorphic information. The training of the Magic Text-Encoder must be conducted after the overall network training is completed. Formally, given a projection $\mathbf{X}\mathbf{W}=\mathbf{Y}$ with $\mathbf{X}\in\mathbb{R}^{q\times w}$, $\mathbf{W}\in\mathbb{R}^{w\times r}$, adapter computes:
$\mathbf{Y}=\mathbf{X}\mathbf{W}+\gamma\mathbf{X}\mathbf{L}_{1}\mathbf{L}_{2},$ (11)
where $\mathbf{X}$ is the input, $\mathbf{Y}$ is the output, $\mathbf{L}_{1}\in\mathbb{R}^{w\times e}$ and $\mathbf{L}_{2}\in\mathbb{R}^{e\times r}$ are adapter parameters, $\mathbf{W}$ is CLIP parameters, $\gamma$ is a scaling coefficient, and $q$, $w$, $e$, $r$ are the dimensions of input, intermediate, low-rank, and output respectively. To avoid losing original capabilities, only $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ are updated.
@@ 4.1 Experimental Settings
Figure 4: Qualitative comparison for metamorphic video generation. Among MagicTime, MakeLongVideo (Make-Long-Video, ), ModelScopeT2v (Modelscope, ), VideoCrafter (Videocrafter, ), Zeroscope (zeroscope, ), LaVie (lavie, ), T2V-Zero (T2V-zero, ), Latte (Latte, ) and Animatediff (animatediff, ), our MagicTime achieves the best results among all T2V models.
@@
Datasets    Metamorphic video generation represents a novel paradigm, yet dedicated evaluation datasets are scarce, presenting challenges for thorough method assessment. To address this need, we utilize the ChronoMagic dataset, the first benchmark specifically designed for metamorphic video generation. Detailed information and examples are provided in the Appendix. This dataset consists of high-quality metamorphic time-lapse videos with corresponding captions. The statistics of the ChronoMagic dataset are presented in Table 2.

### v2

@@ 3.2 ChronoMagic Dataset
where $i$ represents the position of the frame. Ultimately, we segment the video into distinct sections using the identified transition points. As illustrated in Fig. 2a, this process results in a high-quality video dataset with fewer transitions.
Multi-view Text Fusion    After preprocessing, the videos must be annotated with appropriate captions. A simple method involves feeding the video into a multimodal model [58, 59, 60, 61] for video description. However, we observed a significant occurrence of hallucination in open-source models, as described in the Appendix. To address this issue, we adopted In-Context Learning and Chain-of-Thought [62] based on GPT-4V [63, 64] to generate captions in stages. Fig. 2b shows the detailed text generation and fusion process. During this process, we also used titles and hashtags as supplements to reduce model biases. Specifically, we initially employ GPT-4V [63] to generate keyframe captions with the help of titles and hashtags. Subsequently, these captions are used to develop a comprehensive representation of the entire video, culminating in the generation of the final video captions. Additionally, at this stage, GPT-4V evaluates whether the video is a time-lapse and then forms a closed data loop, as described in Section  3.2.
Figure 4: The structure of MagicAdapter and its functional position. MagicAdapter consists of several layers of Linear layers. MagicAdapter-S operates on the output of the Spatial Layer, while MagicAdapter-T functions on the input of the Temporal Layer. $\alpha$ and $\beta$ represent the weight coefficients.
3.3 MagicTime Training Recipe
Magic Adaptive Strategy   The key to encoding metamorphic physical priors into the model lies in decoupling the training process. Video generation models typically consist of spatial layers and temporal layers. The former focuses on learning visually relevant content, while the latter aims to establish motion patterns. To enable the pre-trained text-to-video (T2V) model to generate metamorphic videos, we divide the training process into two phases: spatial training and temporal training, as shown in Fig. 3.
The reasons and benefits are as follows: First, compared to general videos, metamorphic videos contain more static state information (e.g., seeds, leaves, flowers) and dynamic state change information (e.g., seed $\rightarrow$ leaf $\rightarrow$ flower). The decoupled training process can separately encode these two characteristics, thus better encoding new physical priors without losing the original priors. Secondly, the phased training strategy improves the robustness of the model and exhibits better physical consistency when generating videos. Additionally, it reduces training complexity, allowing the model to focus on spatial details and temporal changes separately, thereby improving training stability and efficiency.
To achieve this goal, in the first phase, we remove the temporal layer from the pre-trained model and integrate MagicAdapter-S into the spatial layer. Subsequently, we freeze the remaining model parameters and train using keyframe-text pairs from the ChronoMagic dataset. Since this phase does not involve temporal changes, the model can learn these static features more accurately and also alleviate the impact of video watermarks. In the second phase, we reintroduce the pre-trained temporal layer and propose integrating MagicAdapter-T into the temporal layer. We then freeze other parameters and train using video-text pairs from the ChronoMagic dataset. Since spatial features have been learned in the previous phase, the optimization during the temporal training phase can focus more on the rules of metamorphic changes, thereby avoiding mutual interference between the learning of spatial and temporal features. Our proposed MagicAdapter is simple and lightweight, as shown in Fig. 4. Formally, let $X\in\mathbb{R}^{\text{B}\times\text{C}\times\text{1}\times\text{H}\times\text{W}}$ and $Y\in\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$ be the input to a single spatial layer $\mathcal{F}_{sap}$ and MagicAdapter-S layer $\mathcal{R}_{ada}$, and to a single temporal layer $\mathcal{F}_{temp}$ and MagicAdapter-T layer $\mathcal{M}_{ada}$.
$\quad\mathcal{F}_{outs}(X)=\mathcal{F}_{sap}(X)+\alpha\times\mathcal{R}_{ada}(\mathcal{F}_{sap}(X)),$ (8)
$\quad\mathcal{F}_{outt}(Y)=\mathcal{F}_{temp}(Y)+\beta\times\mathcal{M}_{ada}(Y),$ (9)
where $\mathcal{F}_{outs}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{1}\times\text{H}\times\text{W}}$ and $\mathcal{F}_{outt}\in\mathbb{R}^{\text{B}\times\text{C}\times\text{F}\times\text{H}\times\text{W}}$, $\alpha$ and $\beta$ are scaling coefficients, $B$, $C$, $F$, $H$, and $W$ represent the batch size, channel, frame number, height, and width.
Dynamic Frames Extraction    Open-source algorithms, whether for video generation [65, 66] or multimodal models [58, 67], typically sample a random continuous sequence of $N$ frames for training. However, this approach captures only a brief segment of the full video, leading to limitations such as monotonous actions and restricted variation in amplitude. This is not conducive to extracting training frames for metamorphic videos. By analyzing ChronoMagic’s characteristics, we opt to sample $N$ frames evenly from a video, capitalizing on the nature of metamorphic videos that encompass the entire process of object metamorphosis. This ensures that the training data exhibit metamorphic properties, rather than being confined to general ones, thereby enhancing the generative models’ ability to comprehend physical phenomena. Nonetheless, experiments indicate that uniform sampling from all videos may compromise the model’s original capabilities, resulting in severe flickering phenomena in the generated videos. Therefore, we retain a portion of general videos and dynamically select the sampling strategy based on the results of cascade preprocessing to avoid catastrophic forgetting:
$\mathcal{S}^{*}=\begin{cases}\text{Uniform, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}\leq\delta\\ \text{Random, with }\mathcal{P}\text{, if }\sum_{i=1}^{F-1}\mathcal{T}_{i}>\delta\end{cases}$ (10)
@@ 3.3 MagicTime Training Recipe
TABLE II: Comparison of statistics among Sky Time-lapse [7], Time-lapse-D [8], and the proposed metamorphic ChronoMagic dataset. ChronoMagic is the only open-domain metamorphic time-lapse dataset with video-text pairs.
Dataset ChronoMagic Sky Time-lapse Time-lapse-D
@@
Category Open-domain Sky Sky
Variation Strong (metamorphic) Weak (general) Weak (general)
@@
Resolution 1920 $\times$ 1080 640 $\times$ 360 1920 $\times$ 1080
Information Density Dense Sparse Sparse
Magic Text-Encoder    Textual descriptions corresponding to metamorphic videos typically contain more temporal and state information than those for general videos. Consequently, using a text encoder trained exclusively on general videos is deemed insufficient for this purpose. To overcome this limitation, we introduce the Magic Text-Encoder, which is specifically designed to encode metamorphic prompts while retaining its ability to process general prompts. We select CLIP [57] as our initial text encoder and incorporate several convolution layers to encode metamorphic information while ensuring high efficiency, inspired by Lora [68]. During training, we incorporate the tags ”metamorphic, time-lapse” into the text prompts corresponding to the metamorphic videos. This approach emphasizes the difference between metamorphic and general videos, allowing Magic Text-Encoder to incorporate metamorphic priors without losing its original text-following ability. The training is conducted after the overall network training is completed, because this step is to further refine the model to better interpret the nuances of the metamorphic. Formally, given an original projection $\mathbf{X}_{0}\mathbf{W}_{0}=\mathbf{Y}_{0}$ where $\mathbf{X}_{0}\in\mathbb{R}^{q\times w}$ and $\mathbf{W}_{0}\in\mathbb{R}^{w\times r}$, the Magic Text-Encoder computes:
$\mathbf{Y}_{1}=(\mathbf{X}_{0}+\mathbf{T}_{\text{AG}})\mathbf{W}_{0}+\gamma(\mathbf{X}_{0}+\mathbf{T}_{\text{AG}})\mathbf{L}_{1}\mathbf{L}_{2},$ (11)
where $\mathbf{X}_{0}$ is the input text prompt of CLIP, $\mathbf{Y}_{1}$ is the output of CLIP, $\mathbf{T}_{\text{AG}}$ indicates newly added metamorphic tags, $\mathbf{L}_{1}\in\mathbb{R}^{w\times e}$ and $\mathbf{L}_{2}\in\mathbb{R}^{e\times r}$ are adapter parameters, $\mathbf{W}_{0}$ are the CLIP parameters, $\gamma$ is a scaling coefficient, $q$, $w$, $e$, and $r$ represent the dimensions of input, intermediate, low-rank, and output, respectively. To avoid losing original capabilities, only $\mathbf{L}_{1}$ and $\mathbf{L}_{2}$ are updated. More details of the structure of Magic Text-Encoder can be found in the Appendix.
@@ 4 Experiment
Figure 5: Qualitative comparison for metamorphic video generation. Among MakeLongVideo [69], ModelScopeT2V [5], VideoCrafter2 [70], Zeroscope [71], LaVie [72], T2V-Zero [73], Latte [3] and Animatediff [1], MagicTime achieves best results.
@@ 4.1 Experimental Settings
Datasets    Metamorphic video generation represents a novel paradigm, yet dedicated evaluation datasets are scarce, presenting challenges for thorough method assessment. To address this need, we utilize the ChronoMagic dataset, the first benchmark specifically designed for metamorphic video generation. Detailed information and examples are provided in the Appendix. As shown in Table II, the ChronoMagic dataset is the only metamorphic time-lapse dataset with video-text pairs, which has more categories and stronger variation, that is, higher physical content. Despite the seemingly lower number of videos for ChronoMagic, our overall information density is higher than that of Sky Time-lapse [7] and Time-lapse-D [8], as they mainly consist of general time-lapse videos such as cloud drifting.